OpenStudy (anonymous):

solve the system by the elimination method 5x+3y=-7 7x-2y=13

OpenStudy (anonymous):

Cali, do you want to try these on your own, or no?

OpenStudy (anonymous):

or we can cheat. but we have to do something yes?

OpenStudy (anonymous):

Satellite....lol I'm sure she's taking a test...you don't have to stress her out any more than she already is....lol

OpenStudy (anonymous):

well if i was taking an on line test i would cheat, and not rely on people here giving me goofy answers. i would cheat in such a way that i knew the answers were correct.

OpenStudy (anonymous):

ok, teach me

OpenStudy (anonymous):

Well, the simplest way to do this in my opinion would be to solve both equations for y. If you do that, you can then eliminate y by setting y = y and then solving for x. If you don't know how to solve equations for y, there really isn't much I can do for you.

OpenStudy (bahrom7893):

okay here we go folks!

OpenStudy (anonymous):

the simplest way it here http://www.wolframalpha.com/input/?i=5x%2B3y%3D-7%2C+7x-2y%3D13

OpenStudy (anonymous):

the next simplest way is to use matrices

OpenStudy (anonymous):

Satellite, she said TEACH ME....I guess you just ignored that

OpenStudy (anonymous):

i would never in a million years solve both for y, would you?

OpenStudy (anonymous):

That's the way I've always done it

OpenStudy (bahrom7893):

I was actually trying to teach, but the student is nowhere to be found!

OpenStudy (anonymous):

Solving for y isn't that difficult. I can do it in one step

OpenStudy (anonymous):

multiply equation 1 by 2 and equation 2 by 3 to get \[10x+6y=-14\] \[21x-6y=39\]

OpenStudy (anonymous):

add them upo get \[32x=25\] \[x=\frac{25}{31}\]

OpenStudy (anonymous):

that wasn't too bad was it?

OpenStudy (anonymous):

Sometimes, you can't always solve it that way

OpenStudy (anonymous):

really?

OpenStudy (bahrom7893):

Medals for everyone! and bliss yea you can

OpenStudy (anonymous):

if there is no solution you can't i suppose.

OpenStudy (bahrom7893):

well true..

OpenStudy (anonymous):

Okay...I don't like solving it that way...

OpenStudy (anonymous):

the isolate y way would give u- y=-5/3 x- 7/3 y=3.5x-6.5 I'd go with Satellite's way for this one, that doesn't look terribly fun to add and subtract